likeabossssssss (likeabossssssss):

Identify which table shows a direct variation. Two tables. For each table the column headings are x and y. Table 1 has column 1 values 3, 4.5, 7 and column 2 values 7.5, 11.25, 17.5. Table 2 has column 1 values 1.5, 8, 13 and column 2 values 3, 12, 19.5. A. both tables B. Table 1 only C. Table 2 only D. neither table

likeabossssssss (likeabossssssss):

@antanasia220

OpenStudy (antanasia220):

what do you think the answer is?

likeabossssssss (likeabossssssss):

b

OpenStudy (antanasia220):

idk @ILovePuppiesLol @pooja195

likeabossssssss (likeabossssssss):

@jabez177

likeabossssssss (likeabossssssss):

@IshipDestiel

likeabossssssss (likeabossssssss):

@Hero

hero (hero):

The formula for direct variation is y = kx where k is the constant of variation.

likeabossssssss (likeabossssssss):

im not sure

likeabossssssss (likeabossssssss):

@Hero

hero (hero):

What number can you multiply by 3 to get 7.5?

likeabossssssss (likeabossssssss):

2.5 @Hero

hero (hero):

Okay, check to see if multiplying 4.5 by 2.5 results in 11.25

likeabossssssss (likeabossssssss):

4.5*2.5=11.25

likeabossssssss (likeabossssssss):

@Hero

hero (hero):

Now check the next row in the table.

likeabossssssss (likeabossssssss):

7*2.5=17.5

hero (hero):

So looks like if we divide y by x for Table 1, we get the same k value. What do you think?

hero (hero):

Now see what happens if you try this technique for table 2.

likeabossssssss (likeabossssssss):

1.5*2=3 but u cant get 8* something to get 12

likeabossssssss (likeabossssssss):

@Hero

hero (hero):

For table 1, if you divide y by x, you get a constant k. For table 2, if you divide y by x, you don't get a constant k